# V2XPnP: Vehicle-to-Everything Spatio-Temporal Fusion for Multi-Agent Perception and Prediction

### v2

@@ Abstract
Vehicle-to-everything (V2X) technologies offer a promising paradigm to mitigate the limitations of constrained observability in single-vehicle systems. Prior work primarily focuses on single-frame cooperative perception, which fuses agents’ information across different spatial locations but ignores temporal cues and temporal tasks (e.g., temporal perception and prediction). In this paper, we focus on the spatio-temporal fusion in V2X scenarios and design one-step and multi-step communication strategies (when to transmit) as well as examine their integration with three fusion strategies - early, late, and intermediate (what to transmit), providing comprehensive benchmarks with 11 fusion models (how to fuse). Furthermore, we propose V2XPnP, a novel intermediate fusion framework within one-step communication for end-to-end perception and prediction. Our framework employs a unified Transformer-based architecture to effectively model complex spatio-temporal relationships across multiple agents, frames, and high-definition map. Moreover, we introduce the V2XPnP Sequential Dataset that supports all V2X collaboration modes and addresses the limitations of existing real-world datasets, which are restricted to single-frame or single-mode cooperation. Extensive experiments demonstrate our framework outperforms state-of-the-art methods in both perception and prediction tasks. The codebase and dataset will be released to facilitate future V2X research.
@@ 1 Introduction
Among these strategies, we advocate intermediate fusion within one-step communication, because it effectively balances the trade-off between accuracy and increased transmission load. Moreover, its capability to transmit intermediate spatial-temporal features makes it well-suited for end-to-end perception and prediction, allowing for feature sharing across multiple tasks, as shown in Fig. 1. Based on this strategy, we propose V2XPnP, a V2X spatio-temporal fusion framework leveraging a unified Transformer structure for effective spatial-temporal fusion, encompassing temporal attention, self-spatial attention, multi-agent spatial attention, and map attention. Each agent first extracts its inter-frame and self-spatial features, which can support single-vehicle perception and prediction while reducing the communication load, and then the multi-agent spatial attention model fuses the single-agent feature across different agents.
Another challenge is the lack of real-world sequential datasets encompassing diverse V2X collaboration modes. In V2X scenarios, vehicles and infrastructures serve as primary agents, with collaboration modes that include vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and infrastructure-to-infrastructure (I2I) [19]. Most existing datasets [44, 49] are non-sequential, limited to single collaboration modes, and focus only on single-frame cooperative perception, lacking support for temporal tasks. To bridge this gap, we introduce the first large-scale real-world V2XPnP Sequential Dataset, featuring four agents and supporting all collaboration modes. This dataset includes temporally consistent perception and trajectory data across 100 vehicle-centric (VC) scenarios and 63 infrastructure-centric (IC) scenarios, totaling 40k frames, along with point-cloud and vector maps from 24 collected intersections. The main contributions of this paper can be summarized as follows:
@@
    We present V2XPnP, a V2X spatio-temporal fusion framework with a novel intermediate fusion model within one-step communication. This framework is based on unified Transformer architecture integrating diverse attention fusion modules for V2X spatial-temporal information.
@@ 2 Related Work
V2X Perception and Prediction. V2X perception has been extensively explored, with intermediate fusion emerging as a widely adopted strategy [46, 45, 10]. FFNet [51] and CoBEVFlow [42] utilize historical information (0.5s) from collaborators to mitigate asynchrony, and SCOPE integrates ego-history for detection. However, coordinating multi-agent systems for long-term temporal tasks remains an open challenge. In the prediction domain, deep learning models have been extensively studied for modeling inter-agent interactions [37, 20]. However, the limited short-term visibility of individual vehicles continues to restrict prediction accuracy. Cooperative prediction leveraging V2X has shown potential, though research remains preliminary [53, 41, 48, 34]. To integrate perception and prediction within an end-to-end framework, V2VNet [40] employs a graph neural network for spatio-temporal fusion. UniV2X [52] extended the end-to-end system to support downstream tasks, but with a simplified spatio-temporal fusion module. Despite these advancements, a comprehensive framework for V2X-based spatio-temporal fusion is still lacking.
@@ 3.1 V2X Spatio-Temporal Fusion Strategies
This section addresses what and when to transmit for spatio-temporal fusion, and Sec. 3.2 delves into the fusion process.
What to Transmit. In multi-agent spatial fusion, three fusion strategies are widely adopted in single-frame cooperative perception, i.e., early, late, and intermediate fusion [15, 49], which involve the transmission of raw perception data, bounding boxes, and intermediate features, respectively. Adopting this framework, we extend these fusion approaches to the V2X spatio-temporal fusion context, as illustrated in Fig. 2. (1) Early fusion transmits the entire raw historical perception data to retain complete feature information but imposes the highest transmission load. (2) Late fusion shares only the final detected results at each historical frame, resulting in the lowest transmission load but losing most of the feature information. (3) Intermediate fusion transmits intermediate spatio-temporal BEV features, striking a balance between information quality and transmission load.
@@ 3.2 V2XPnP Framework
The spatio-temporal features of intermediate fusion render it a natural fit for end-to-end perception and prediction. Accordingly, we propose a unified end-to-end perception and prediction framework to perform multiple tasks across spatio-temporal dimensions. The overall V2XPnP framework is illustrated in Fig. 3, which includes six components and is unfolded in this section. The detail of the spatio-temporal fusion model is provided in Sec. 3.3. Notably, each module in V2XPnP is modular, allowing for easy replacement.
@@
Detection and Prediction Heads. Finally, a detection head and a prediction head are connected to the final feature $\mathbf{F}$ to output states for each predefined anchor box: (1) The detection head contains two 2D convolution layers for bounding box regression and classification. The regression branch outputs $(x,y,z,w,l,h,\theta)$, which represents the bounding box position, size, and yaw. The classification branch outputs the confidence score of each anchor box, determining whether it corresponds to an object or background. (2) The prediction head outputs offset values for each anchor box at each future timestamp using two 2D convolution layers, and the final trajectory is generated by accumulating these offsets.
@@ 3.3 Spatio-Temporal Fusion Transformer
In this section, we introduce spatio-temporal fusion with a unified Transformer architecture. The proposed model comprises three blocks: temporal attention, self-spatial attention, and multi-agent spatial attention, as shown in Fig. 3, and two core fusion modules. (1) Multi-frame temporal fusion: Each agent first extracts their spatio-temporal features through iterative temporal and self-spatial attentions. (2) Multi-agent spatial fusion: rich BEV features from multiple agents are acquired via V2X and then fused through iterative multi-agent spatial and self-spatial attentions.
@@ 3.4 Learning Objective
The learning objective is comprised of temporal perception and prediction tasks. First, we define the perception loss as the combination of regression $\mathcal{L}_{\text{reg}}$ and classification loss $\mathcal{L}_{\text{cla}}$ of the predefined anchor box. Specifically, the smooth $\ell_{1}$-loss is leveraged for the regression part, and the focal loss [31] is utilized for classification. Second, we define the prediction loss $\mathcal{L}_{\text{pred }}$ as $\ell_{2}$-loss between the prediction points sequence with ground truth trajectory. The final loss function is the weighted sum of $\mathcal{L}_{\text{reg }},\mathcal{L}_{\text{cla }}$, and $\mathcal{L}_{\text{pred }}$.
@@ 4 V2XPnP Sequential Dataset
We introduce the V2XPnP-Sequential dataset, the first large-scale, real-world V2X sequential dataset featuring multiple agents and all collaboration modes. This dataset comprises 100 scenarios (49 2V+2I scenarios, 42 V2V scenarios, and 9 V+2I scenarios), each spanning 95 to 283 frames captured at 10 Hz. The dataset comprises two data sequences from CAV perception (point clouds and camera images) and two data sequences from infrastructure perception, as shown in Fig. 4(b). We also provide corresponding vector maps and point-cloud maps for all collection areas, as shown in Fig. 4(c). Ten object categories are included, and the average trajectory length and frequency of each category are shown in Fig. 4(d). Further details on data visualization, annotation, trajectory and map generation are provided in the supplementary.
@@ 4.1 Data Acquisition
V2X temporal tasks require diverse time-consistent perception data and object behavior data. We choose urban arterial roads, expressways, and intersections to collect V2X data. The sensor configurations for two CAVs and two infrastructures are shown in Fig. 4(a). By permuting and combining the behavior patterns of CAVs (such as overtaking, platooning, turning, etc.), we designed a total of about 60 interaction pairs between CAVs to collect data. From more than 66h driving logs, we annotate 32 representative scenarios. We also process the non-sequential data from V2X-Real [44] with a V2X sequential data processing pipeline, detailed in Sec. 4.2. Since perception tasks do not need to consider labeling consistency, identical objects may be assigned different IDs within a sequence, leading to data fragmentation and limiting support for temporal tasks. The final dataset is yielded by processing our collected data and V2X-Real data.
@@ 5.1 Experimental Setup
$\text{EPA}=\frac{|\hat{S}_{\text{match, hit}}|-\alpha N_{\text{FP}}}{N_{\text{GT}}},$ (4)
@@
Implementation Details. During the testing stage, we select a fixed agent as the ego agent in each cooperative scenario, while the ego agent is shuffled and randomly selected during training. Following the real-time setting [55], we set the communication range to 50 meters and evaluate surrounding agents within a range of $x\in[-70,70]$m and $y\in[-40,40]$m. Messages beyond 50 meters are discarded. Besides, the history length is 2s (2Hz), and the prediction horizon is 3s (2Hz). The train/validation/test data splits are 80/6/14 scenarios. Additional training and model details are provided in the supplementary materials.
@@ 5.3 Results
Tab. 1 presents the benchmark results across four V2X collaboration modes. Since prediction performance inherently depends on detection quality, the difficulty posed by different detected objects can significantly impact prediction accuracy. Thus, the EPA metric serves as the most appropriate indicator for assessing the overall performance. More details and baseline results are provided in the supplementary.
@@
When to Transmit. Tab. 2 shows the performance of our model under two communication strategies. We find that the perception and prediction performance of one-step communication is improved compared to multi-step communication by 5.0% AP and 12% in EPA. This improvement arises because each agent first directly fuses their lossless raw temporal data before sharing, avoiding error accumulation from lossy intermediate information transformed in multi-step communication across temporal dimensions. Moreover, the one-step strategy compresses spatio-temporal feature transmission (under a 32$\times$ compression rate) from 5$\times$0.269 Mb to 0.269 Mb compared to multi-step communication, while mitigating information loss when agents move out of communication range in historical frames. At a typical C-V2X data transmission rate [4], the transmission delay of one-step communication is approximately $10\sim 20$ ms.
How to Fuse. Tab. 3 provides the ablation study of V2XPnP, showing the effectiveness of different components in V2XPnP. The temporal fusion module provides the history information for current-frame detection, while multi-agent spatial fusion alleviates occlusions and improves performance by incorporating other views. The map fusion module enhances trajectory prediction by guiding future trajectories to align with road structures. Our complete V2XPnP model with all these fusion modules performs the best.
@@
Infrastructure vs. Vehicle Centric. As shown in Tab. 1, models under VC and IC modes outperform those in V2V and I2I modes, because VC and IC can aggregate information from up to four agents rather than only two, resulting in enhanced environmental understanding. Notably, the evaluation protocol is consistent across all models within each mode. Additionally, stationary infrastructure-based agents in IC and I2I modes offer higher prediction accuracy by providing elevated sensing perspectives and less noisy data.
Transmission Data Size and Robustness Test. Results in Fig. 6 indicate that V2XPnP achieves a good balance between communication efficiency and accuracy, maintaining superior performance compared to full-size V2X-ViT* even at a ×$\times 128$ compression rate. Following the setting in [45], we provide the results with 100-500 ms time delay and positional/head Gaussian noise from (0.2m, 0.2°) to (1m, 1°). Both V2XPnP and V2X-ViT* maintain robust performance due to their spatial attention fusion, and V2XPnP performs better due to designed temporal attention.
Qualitative Results. Fig. 5 visualizes the outcomes of cooperative perception and prediction across different fusion models. The No Fusion model is constrained by its limited field of view. The FaF model, leveraging temporal information within an end-to-end pipeline, performs better under occlusion. Late and early fusion models significantly benefit from multi-agent data integration, though late fusion remains impacted by error propagation, such as detection heading errors misleading trajectory direction. Notably, the end-to-end intermediate fusion model, particularly V2XPnP, performs better in both detection and prediction tasks.
@@ A.2 V2XPnP Model Details
Multi-agent Spatial Attention. Our dataset categorizes agents as infrastructure agents, denoted by negative labels (i.e., $-1$ and $-2$), or connected automated vehicles (CAV) agents, denoted by positive labels (i.e., $1$ and $2$). To capture the heterogeneous dependencies among these agents, we construct a heterogeneous graph and employ distinct attention fusion parameters for each agent type. The multi-agent spatial attention utilizes eight attention heads, and we stack three multi-agent spatial fusion modules with the multi-agent spatial attention to capture the inter-agent relationships.
@@ A.3 Loss Function
This section provides the loss function employed in our multi-task model. The initial weights of regression loss $\mathcal{L}_{\text{reg }}$, classification loss $\mathcal{L}_{\text{cla}}$ and prediction loss $\mathcal{L}_{\text{pred}}$ are set as $w_{\text{reg}},w_{\text{cla}},w_{\text{pred}}=2.0,1.0,2.0$. For single-task learning, the same loss function is used but weights exclusively on the components relevant to that task.
@@ B Additional Benchmark Results
We present additional benchmark results for V2VNet [40], F-Cooper [6], and DiscoNet [27] across all collaboration modes, as shown in Tab. S1. Our proposed V2XPnP consistently outperforms these SOAT baselines in terms of EPA and AP across all collaboration modes. Notably, V2VNet* exhibits lower performance, likely due to the absence of a map and the loss of temporal features during explicit feature ROI matching.

### v3

@@ Abstract
Vehicle-to-everything (V2X) technologies offer a promising paradigm to mitigate the limitations of constrained observability in single-vehicle systems. Prior work primarily focuses on single-frame cooperative perception, which fuses agents’ information across different spatial locations but ignores temporal cues and temporal tasks (e.g., temporal perception and prediction). In this paper, we focus on the spatio-temporal fusion in V2X scenarios and design one-step and multi-step communication strategies (when to transmit) as well as examine their integration with three fusion strategies - early, late, and intermediate (what to transmit), providing comprehensive benchmarks with 11 fusion models (how to fuse). Furthermore, we propose V2XPnP, a novel intermediate fusion framework within one-step communication for end-to-end perception and prediction. Our framework employs a unified Transformer-based architecture to effectively model complex spatio-temporal relationships across multiple agents, frames, and high-definition maps. Moreover, we introduce the V2XPnP Sequential Dataset that supports all V2X collaboration modes and addresses the limitations of existing real-world datasets, which are restricted to single-frame or single-mode cooperation. Extensive experiments demonstrate that our framework outperforms state-of-the-art methods in both perception and prediction tasks.
@@ 1 Introduction
Among these strategies, we advocate intermediate fusion within one-step communication because it effectively balances the trade-off between accuracy and increased transmission load. Moreover, its capability to transmit intermediate spatial-temporal features makes it well-suited for end-to-end perception and prediction, allowing for feature sharing across multiple tasks, as shown in Fig.˜1. Based on this strategy, we propose V2XPnP, a V2X spatio-temporal fusion framework leveraging a unified Transformer structure for effective spatial-temporal fusion, encompassing temporal attention, self-spatial attention, multi-agent spatial attention, and map attention. Each agent first extracts its inter-frame and self-spatial features, which can support single-vehicle perception and prediction while reducing the communication load, and then the multi-agent spatial attention model fuses the single-agent features across different agents.
Another challenge is the lack of real-world sequential datasets encompassing diverse V2X collaboration modes. In V2X scenarios, vehicles and infrastructures serve as primary agents, with collaboration modes that include vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and infrastructure-to-infrastructure (I2I), vehicle-centric (VC), and infrastructure-centric (IC) [19]. Most existing datasets [44, 49] are non-sequential, limited to single collaboration modes, and focus only on single-frame cooperative perception, lacking support for temporal tasks. To bridge this gap, we introduce the first large-scale real-world V2XPnP Sequential Dataset, featuring four agents and supporting all collaboration modes. This dataset includes temporally consistent perception and trajectory data across 49 2V+2I scenarios, 38 V2V scenarios, and 9 V+2I scenarios, totaling 40k frames, along with point-cloud and vector maps from 24 collected intersections. The main contributions of this paper can be summarized as follows:
@@
    We present V2XPnP, a V2X spatio-temporal fusion framework with a novel intermediate fusion model within one-step communication. This framework is based on a unified Transformer architecture integrating diverse attention fusion modules for V2X spatial-temporal information.
@@ 2 Related Work
V2X Perception and Prediction. V2X perception has been extensively explored, with intermediate fusion emerging as a widely adopted strategy [46, 45, 10]. FFNet [51] and CoBEVFlow [42] utilize historical information (0.5s) from collaborators to mitigate asynchrony, and SCOPE integrates ego-history for detection. However, coordinating multi-agent systems for long-term temporal tasks remains an open challenge. In the prediction domain, deep learning models have been extensively studied for modeling inter-agent interactions [37, 20]. However, the limited short-term visibility of individual vehicles continues to restrict prediction accuracy. Cooperative prediction leveraging V2X has shown potential, though research remains preliminary [53, 41, 48, 34].
To integrate perception and prediction within an end-to-end framework, V2VNet [40] employs a graph neural network for spatio-temporal fusion. UniV2X [52] extended the end-to-end system to support downstream tasks, but with a simplified spatio-temporal fusion module. Despite these advancements, a comprehensive framework for V2X-based spatio-temporal fusion is still lacking.
@@ 3.1 V2X Spatio-Temporal Fusion Strategies
This section addresses what and when to transmit for spatio-temporal fusion, and Sec.˜3.2 delves into the fusion process.
What to Transmit. In multi-agent spatial fusion, three fusion strategies are widely adopted in single-frame cooperative perception, i.e., early, late, and intermediate fusion [15, 49], which involve the transmission of raw perception data, bounding boxes, and intermediate features, respectively. Adopting this framework, we extend these fusion approaches to the V2X spatio-temporal fusion context, as illustrated in Fig.˜2. (1) Early fusion transmits the entire raw historical perception data to retain complete feature information but imposes the highest transmission load. (2) Late fusion shares only the final detected results at each historical frame, resulting in the lowest transmission load but losing most of the feature information. (3) Intermediate fusion transmits intermediate spatio-temporal BEV features, striking a balance between information quality and transmission load.
@@ 3.2 V2XPnP Framework
The spatio-temporal features of intermediate fusion render it a natural fit for end-to-end perception and prediction. Accordingly, we propose a unified end-to-end perception and prediction framework to perform multiple tasks across spatio-temporal dimensions. The overall V2XPnP framework is illustrated in Fig.˜3, which includes six components and is unfolded in this section. The detail of the spatio-temporal fusion model is provided in Sec.˜3.3. Notably, each module in V2XPnP is modular, allowing for easy replacement.
@@
Detection and Prediction Heads. Finally, a detection head and a prediction head are connected to the final feature $\mathbf{F}$ to output states for each predefined anchor box: (1) The detection head contains two 2D convolution layers for bounding box regression and classification. The regression branch outputs the bounding box position $(x,y,z)$, size $(w,l,h)$, and yaw $\theta$. The classification branch outputs the confidence score of each anchor box, determining whether it corresponds to an object or background. (2) The prediction head outputs offset values for each anchor box at each future timestamp using two 2D convolution layers, and the final trajectory is generated by accumulating these offsets.
@@ 3.3 Spatio-Temporal Fusion Transformer
In this section, we introduce spatio-temporal fusion with a unified Transformer architecture. The proposed model comprises three blocks: temporal attention, self-spatial attention, and multi-agent spatial attention, as shown in Fig.˜3, and two core fusion modules. (1) Multi-frame temporal fusion: Each agent first extracts their spatio-temporal features through iterative temporal and self-spatial attentions. (2) Multi-agent spatial fusion: rich BEV features from multiple agents are acquired via V2X and then fused through iterative multi-agent spatial and self-spatial attentions.
@@ 3.4 Learning Objective
[…]
@@ 4 V2XPnP Sequential Dataset
We introduce the V2XPnP-Sequential dataset, the first large-scale, real-world V2X sequential dataset featuring multiple agents and all collaboration modes. This dataset comprises 96 scenarios (49 2V+2I scenarios, 38 V2V scenarios, and 9 V+2I scenarios), each spanning 95 to 283 frames captured at 10 Hz. The dataset comprises two data sequences from CAV perception (point clouds and camera images) and two data sequences from infrastructure perception, as shown in Fig.˜4(b). We also provide corresponding vector maps and point-cloud maps for all collection areas, as shown in Fig.˜4(c). Ten object categories are included, and the average trajectory length and frequency of each category are shown in Fig.˜4(d). Further details on data visualization, annotation, trajectory and map generation are provided in the supplementary.
@@ 4.1 Data Acquisition
V2X temporal tasks require diverse time-consistent perception data and object behavior data. We choose urban arterial roads, expressways, and intersections to collect V2X data. The sensor configurations for two CAVs and two infrastructures are shown in Fig.˜4(a). By permuting and combining the behavior patterns of CAVs (such as overtaking, platooning, turning, etc.), we designed a total of about 60 interaction pairs between CAVs to collect data. From more than 66h driving logs, we annotate 28 representative scenarios. We also process the non-sequential data from V2X-Real [44] with a V2X sequential data processing pipeline, detailed in Sec.˜4.2. Since perception tasks do not need to consider labeling consistency, identical objects may be assigned different IDs within a sequence, leading to data fragmentation and limiting support for temporal tasks. The final dataset is yielded by processing our collected data and V2X-Real data.
@@ 5.1 Experimental Setup
$\vskip-8.5359pt\text{EPA}=\frac{|\hat{S}_{\text{match, hit}}|-\alpha N_{\text{FP}}}{N_{\text{GT}}},\vskip-2.84544pt$ (4)
@@
Implementation Details. During the testing stage, we select a fixed agent as the ego agent in each cooperative scenario, while the ego agent is shuffled and randomly selected during training. Following the real-time setting [55], we set the communication range to 50 meters and evaluate surrounding agents within a range of $x\in[-70,70]$m and $y\in[-40,40]$m. Messages beyond 50 meters are discarded. Besides, the history length is 2 s (2 Hz), and the prediction horizon is 3 s (2 Hz). The train/validation/test data splits are 76/6/14 scenarios. Additional training and model details are provided in the supplementary materials.
@@ 5.3 Results
Tab.˜1 presents the benchmark results across four V2X collaboration modes. Since prediction performance inherently depends on detection quality, the difficulty posed by different detected objects can significantly impact prediction accuracy. Thus, the EPA metric serves as the most appropriate indicator for assessing the overall performance. More details and baseline results are provided in the supplementary.
@@
[…]
How to Fuse. Tab.˜3 provides the ablation study of V2XPnP, showing the effectiveness of different components in V2XPnP. The temporal fusion module provides the history information for current-frame detection, while multi-agent spatial fusion alleviates occlusions and improves performance by incorporating other views. The map fusion module enhances trajectory prediction by guiding future trajectories to align with road structures. Our complete V2XPnP model with all these fusion modules performs the best.
@@
Infrastructure vs. Vehicle Centric. As shown in Tab.˜1, models under VC and IC modes outperform those in V2V and I2I modes, because VC and IC can aggregate information from up to four agents rather than only two, resulting in enhanced environmental understanding. Notably, the evaluation protocol is consistent across all models within each mode. Additionally, stationary infrastructure-based agents in IC and I2I modes offer higher prediction accuracy by providing elevated sensing perspectives and less noisy data.
Transmission Data Size and Robustness Test. Results in Fig.˜6 indicate that V2XPnP achieves a good balance between communication efficiency and accuracy, maintaining superior performance compared to full-size V2X-ViT* even at a $\times 128$ compression rate. Following the setting in [45], we provide the results with 100-500 ms time delay and positional/head Gaussian noise from (0.2m, 0.2°) to (1m, 1°). Both V2XPnP and V2X-ViT* maintain robust performance due to their spatial attention fusion, and V2XPnP performs better due to its designed temporal attention.
Qualitative Results. Fig.˜5 visualizes the outcomes of cooperative perception and prediction across different fusion models. The No Fusion model is constrained by its limited field of view. The FaF model, leveraging temporal information within an end-to-end pipeline, performs better under occlusion. Late and early fusion models significantly benefit from multi-agent data integration, though late fusion remains impacted by error propagation, such as detection heading errors misleading trajectory direction. Notably, the end-to-end intermediate fusion model, particularly V2XPnP, performs better in both detection and prediction tasks.
@@ 6 Conclusions
Acknowledgements
This work was supported by the Federal Highway Administration Center of Excellence on New Mobility and Automated Vehicles, and by the National Science Foundation under Award No. 2346267, POSE: Phase II - DriveX: An Open Source Ecosystem for Automated Driving and Intelligent Transportation Research.
@@ A.2 V2XPnP Model Details
Multi-agent Spatial Attention. Our dataset categorizes agents as infrastructure agents, denoted by negative labels (i.e., 1-1 and 2-2), or connected automated vehicles (CAV) agents, denoted by positive labels (i.e., $1$ and $2$). To capture the heterogeneous dependencies among these agents, we construct a heterogeneous graph and employ distinct attention fusion parameters for each agent type. The multi-agent spatial attention utilizes eight attention heads, and we stack three multi-agent spatial fusion modules with the multi-agent spatial attention to capture the inter-agent relationships.
@@ A.3 Loss Function
[…]
@@ B Additional Benchmark Results
We present additional benchmark results for V2VNet [40], F-Cooper [6], and DiscoNet [27] across all collaboration modes, as shown in Tab.˜S1. Our proposed V2XPnP consistently outperforms these SOAT baselines in terms of EPA and AP across all collaboration modes. Notably, V2VNet* exhibits lower performance, likely due to the absence of a map and the loss of temporal features during explicit feature ROI matching.